2) Use Laplace Transforms to solve the initial value problem y/ -3y = -10(1-2) with y(0) = 8/9.
Answer #1

y-3y=t-40(+-2), ylor Taking Laplace transform on both sides, we get L{y} - 32{y} = 2{t} - 4.2{U (+-2)} sÝ(s) = y(0) – 3 Y(5)